OpenStudy (faman39):

After how many months will Rita pay $200 in simple annual interest on a credit card that charges 12% on her $400 balance?

OpenStudy (dumbcow):

12% is per year right?

OpenStudy (dumbcow):

wait what is the monthly payment

OpenStudy (faman39):

it not given there

OpenStudy (anonymous):

i think you gotta use a exponential function dude, look in your notes and post it here.

OpenStudy (anonymous):

I = Prt I= interest in dollars P = the amount of money r = the rate t = time 200 = (400)*(0.12)(t) t = 200/(400*0.12) This is just a guess, it's been a long time since I've had to do these problems :-/

OpenStudy (anonymous):

And I think time is in years.

OpenStudy (anonymous):

Hi Faman! Long time!

OpenStudy (faman39):

hey data!

OpenStudy (anonymous):

cuz its gona be 400 times .12 for the first year so the total will be 448 aftier the first year

OpenStudy (dumbcow):

ok i'll assume you are only paying the interest each month 400*r*t = 200 monthly rate is 1% 4 t = 200 t = 50 months

OpenStudy (anonymous):

It seems you're getting a lot of help with this. Its nice to see you again!

OpenStudy (faman39):

so how r u doin? its been a long time, right!

OpenStudy (anonymous):

I'm just fine. Loosing some sleep for the sake of math =)

OpenStudy (faman39):

i missed u guys lot, thing is, i was little busy nowaday, tats y i cant come here like b4. :(

OpenStudy (anonymous):

NM! I didn't REEEAD the problem :-P

OpenStudy (faman39):

ok no prob lol!

OpenStudy (anonymous):

Okay when you solve for t, it's going to be in years. Convert your years into months.

OpenStudy (faman39):

u still need to take care of urself, data.

OpenStudy (faman39):

do u know tat?

OpenStudy (anonymous):

Lol he left

OpenStudy (faman39):

lol i still can see him here, weird.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

I think he was helping someone factor

OpenStudy (faman39):

no brit told u left, ur here data!

OpenStudy (anonymous):

@faman39 I got 50 months for your problem.

OpenStudy (faman39):

lol!

OpenStudy (anonymous):

but you said I need to take care of myself? I did not understand that

OpenStudy (anonymous):

I think she means don't kill yourself over math... oh boy you're gonna have fun in multivariable calculus!

OpenStudy (faman39):

no ur sayin tat ur losin ur sleep for math, tats y .

OpenStudy (anonymous):

ahh ok!

OpenStudy (faman39):

ya brit is rite!

OpenStudy (anonymous):

pls explain to mee brine!

OpenStudy (faman39):

thank u so much for medals guys.

OpenStudy (faman39):

oh ur asking from brit or me?

OpenStudy (anonymous):

I gave him a triple integral problem.

OpenStudy (faman39):

Lol

OpenStudy (anonymous):

The purpose of them is for finding the volume of a weird shape. http://tutorial.math.lamar.edu/Classes/CalcIII/TripleIntegrals.aspx

OpenStudy (anonymous):

Well the region of integration is a volume.

OpenStudy (anonymous):

Yeah mr. technical

OpenStudy (faman39):

lol

OpenStudy (anonymous):

hehe sorry =P

OpenStudy (anonymous):

mr. snooty math engineer person

OpenStudy (anonymous):

lolz

OpenStudy (anonymous):

what is the meaning of snooty?

OpenStudy (anonymous):

uh... pretentious is a close word? j/k though

OpenStudy (anonymous):

sorry =(

OpenStudy (anonymous):

I'm kidding

OpenStudy (anonymous):

@faman39 you'd better make us proud and take calculus!

OpenStudy (anonymous):

Ok. I'm going to sleep I'm loosing my mind with this problem.

OpenStudy (faman39):

lol me.

OpenStudy (faman39):

no i m gonna be worst

OpenStudy (anonymous):

Omg don't let multi get to you yet! You'd not even in the class LOL Hey I should send you some of the old exams my instructor gave us. I have NO idea how I passed.

OpenStudy (anonymous):

*you're

OpenStudy (anonymous):

@faman39 there are so many resources out there. Like Paul's online notes and TheIntegralCalc's channel. That girl is the best at explaining things.

OpenStudy (faman39):

But ur really brilliant tats y, without knowin anything, lol.

OpenStudy (faman39):

u passed ur exams.

OpenStudy (anonymous):

t = 200 * 12 / 48 = 50 mos

OpenStudy (faman39):

Hi chloro.

OpenStudy (faman39):

whats up?

OpenStudy (anonymous):

::)

OpenStudy (faman39):

:)